OpenStudy (anonymous):

I made a mistake but im not sure what it is. What are the foci of the ellipse given by the equation (x-2)^2/ 36 + y-8^2/144 i did h =+/- a k 2 +/- c 8 c^2 = 64 - 4 c='s sqrt 60 but my options are 2 (+/-) 6 sqrt 3, 8 2 (+/-) 6 sqrt 5 , 8 2, 8(+/-) 6 sqrt 3 2,8 (+/-) 6 sqrt 5

OpenStudy (anonymous):

third option ..

OpenStudy (anonymous):

since b>a, the ellipse is gonna have a major axis horizontally.. c^2=b^2-b^2 c=+-sqrt(144-36)=+-sqrt(108) the center is at (2,8) .: the coordinate of foci are gonna be (2,8+-sqrt(108)

OpenStudy (anonymous):

vertically *

OpenStudy (anonymous):

cann you tell me what my mistake was?

OpenStudy (anonymous):

um sorry im not sure about your method ..

OpenStudy (anonymous):

ok thanks anyway